# Monte Carlo Tree Diffusion for System 2 Planning

### v3

@@ A.3 Model Hyperparameters
For reproducibility, we provide details on the hyperparameters used in our experiments. These settings were chosen based on prior work and empirical tuning to ensure stable training and evaluation across all tasks. The almost same hyperparameters were applied consistently across tasks, except in cases where task-specific configurations are required. They will be discussed in each task section.
@@ A.3.1 Diffuser
We build our tested Diffuser on the official implementation, which is available https://github.com/jannerm/diffuser. While keeping the core architecture and training procedure intact, we made the following modifications to align with our experimental setup:
  • Guidance Function: For pointmaze tasks, Diffuser originally applied the goal-inpainting guidance, but in our evaluation, we employed a distance-based reward function to ensure fair comparison across baselines.
  • Replanning Variant: To evaluate the impact of test-time iterative refinement, we implemented a replanning strategy that periodically re-samples trajectories based on the agent’s updated state.
  • Random Search Variant: To evaluate the impact of test-time compute scaling with more samples, we implemented Diffuser version ”Sample-Score-Rank” (d-mpc). Additionally, each sample can be guided through different guidances, we applied this by uniformly distributing the guidance set $[0.01,0.05,0.1,0.2,0.3]$ to the samples.
@@ A.3.2 Diffusion Forcing
We build our tested Diffusion Forcing upon the officially released code from https://github.com/buoyancy99/diffusion-forcing. The following modifications were made:
  • Replanning Frequency: Since Diffusion Forcing inherently supports test-time replanning, we evaluated it under comparable conditions by applying periodic plan refinement, similar to Diffuser-Replanning.
  • Transformer Backbone: To maintain consistency across models, we utilized the Transformer-based backbone provided in the Diffusion Forcing repository, rather than alternative architectures.
@@ A.3.3 Monte Carlo Tree Diffusion (MCTD)
MCTD follows Diffusion Forcing hyperparameters, but some of them are modified for MCTD and tree search related hyperparameters are added.
@@ A.3.4 Value-Learning Policy
For some tasks, to handle complex action space, we applied a value-learning policy by modifying the PlanDQ (plandq) source code, https://github.com/changchencc/plandq.
@@ A.4 The Evaluation Details
We followed the OGBench (4_park2024ogbench) task configurations, 5 tasks per each environment. For each task, we evaluated 10 random seeds per each model, and reported the averaged success rates and their standard deviations.
A.5 Maze Navigation with Point‐Mass and Ant Robots
A.5.1 Modifications from OGBench
Although the original OGBench datasets include random noise in start and goal positions, we remove this noise to isolate model performance from environmental randomness. This modification facilitates clearer comparisons across methods and ensure that differences in performance can be attributed to planning capabilities rather than uncontrolled stochasticity. Additionally, we incorporate velocity into the state representation, which the original benchmark does not provide, in order to apply the heuristic controller for pointmaze tasks implemented in (diffuser). For antmaze giant task, we found there are some cases where the plans are well generated, but due to the inaccurate execution, it cannot be succeed. To prevent this, we extend the episode length from $1000$ to $1500$.
A.5.2 Guidance Function
To plan long-horizon while sustaining the plan effectiveness to reach the goal as fast as possible, we applied the guidance function to reduce the distances between every state and goal, $\sum_{i}||x_{i}-g||_{2}$ (diffusion_forcing). We note that, to do fair comparison we applied the same guidance style over every baseline. For example, Diffuser originally used the goal-inpainting guidance for pointmaze, but in this evaluation, we used the distance guidance.
A.5.3 Guidance Set
We applied different guidance sets to MCTD for the different tasks. For pointmaze medium and large, $[0,0.1,0.5,1,2]$ is applied, and for pointmaze giant, $[0.5,1,2,3,4]$ is applied. For antmaze tasks, $[0,1,2,3,4,5]$ is set.
A.5.4 Planning Horizon
For medium and large mazes, we applied $500$ planning horizons and $1000$ for giant to train the models with enough data. Medium and large navigate dataset trajectory horizons are $1000$ and $2000$ for giant. To make more data through sliding window technique, we applied smaller planning horizon than the dataset horizon.
A.5.5 Heuristic Controller and Value-Learning Policy
For pointmaze, we take the heuristic controller designed in (diffuser). For antmaze, we use the value-learning policy (dql) as done in (plandq). However, we note that different from (plandq), we only give the $10$ steps further planned position information as the subgoal for that.
A.5.6 MCTD Reward Function
For the reward function, we designed heuristic rules. One of those is to check whether there exists too large position difference between near states, because it is impossible physically. Another rule is to give a reward when reaching the goal. Earlier reaching can get the larger reward, $r=(H-t)/H$, where $H$ and $t$ are the horizon length and current step.
@@ A.5.6 MCTD Reward Function
A.5.7 Short‐Horizon (Stitch) Results
To examine performance when only short-horizon trajectories are given in the training phase, we evaluate on the stitch datasets, summarized in Table 8. The trajectory horizon in stitch datasets is much shorter than the required steps to reach the goal (e.g., the horizon in the dataset is 100 while over 200 steps are required in medium map tasks), we applied replanning on Diffuser, Diffusion Forcing, and MCTD. Different from the long-horizon dataset given cases, the Diffuser shows better performances than Diffusion Forcing, it can be seen that the implicit stitching ability of the Diffuser is better than Diffusion Forcing. On the other hand, MCTD shows better performance than baselines in the medium-sized map, due to multiple samplings with exploration, but the search length could not be long due to the limited length of trajectories in the training dataset, so it does not show good performance in large map.
Table 9: Comparison Results with Goal‐Inpainting Diffuser. Success rates (%) on pointmaze and antmaze environments with medium, large, and giant mazes for the navigate datasets. Parenthetical values in the AntMaze Giant row denote performance when interacting with more steps (1500 steps) than predefined steps in OGBench (1000 steps).
@@ A.5.7 Short‐Horizon (Stitch) Results
giant-navigate-v0 $0\pm 0$ ($0\pm 0$) $0\pm 0$ ($0\pm 0$) $12\pm 10$ $20\pm 0$ $50\pm 16$ ($\textbf{94}\pm\textbf{9}$)
A.5.8 Comparison with Goal‐Inpainting Diffuser
@@
Diffuser (diffuser) typically employs a learned guidance function to bias generated trajectories toward high‐return outcomes. For pointmaze tasks, however, an alternative approach—goal‐inpainting guidance—has also been introduced (diffuser), which further improves single‐pass planning by “imagining” an intermediate trajectory whose final state is the designated goal. In the context of Diffuser‐Replanning, we adapt goal‐inpainting by providing the experienced trajectory as contextual information rather than a direct target for denoising.
@@
As shown in Table 9, goal‐inpainting outperforms standard Diffuser with heuristic guidance across medium, large, and giant pointmazes. By treating the goal as the terminal state to inpaint, the Diffuser model effectively samples trajectories from both ends, leveraging local convolutional mechanisms to reconcile the start and goal conditions. Although this approach can yield marked performance gains, MCTD remains superior in long‐horizon scenarios, outperforming both the base and replanning variants of goal‐inpainting Diffuser. The key distinction lies in MCTD’s tree‐structured partial denoising and adaptive branching: while goal inpainting enhances single‐pass generation, it lacks the iterative exploration‐exploitation loop of explicit tree search. Moreover, the advantage of two‐ended trajectory imagination diminishes when the maze is extremely large (e.g., giant mazes). Consequently, MCTD uncovers robust solutions even where goal inpainting begins to show diminishing returns, underscoring the benefits of tree‐based refinement in complex, long‐range planning tasks.
A.6 Robot Arm Cube Manipulation
A.6.1 Object‐Wise Guidance
@@
Because the robot arm can only move a single cube at any given time, we adopt an object‐wise guidance mechanism in MCTD. Rather than applying a single holistic diffusion schedule to all cubes simultaneously, MCTD treats the selection of which cube to move next as a meta‐action. This design partially alleviates issues that arise when a single plan attempts simultaneous motion of multiple cubes, though multi‐object manipulation still demands careful sequencing to avoid suboptimal interleavings.
A.6.2 Guidance Function
@@
The guidance function itself is same to that for maze tasks, Appendix A.5.2, but the object-specific guidance is applied to MCTD and MCTD-Replanning. The guidance tries to reduce the distances between specific object states and the object’s goal only.
A.6.3 Guidance Set
@@
For cube tasks, the guidance set $[1,2,4]$ for each object is applied. For example, for two cube tasks, the size of the guidance set is $6$ with three guidance set for each object.
A.6.4 Reward Function for Robot Arm Cube Manipulation
In addition to the reward function for the maze environment, we added some rules not to generate unrealistic plans. One of them is not to move multiple objects simultaneously. The small noise on planning is okay, but if multiple objects move adequately, the reward function decides the plan is unrealistic and gives a 00 reward. Other rules are, that the final state of the plan cannot be in the air, it should be on top of another cube or on the floor. The cube cannot posit where another cube is located, and when there is another cube on top of the target cube, the plan to move that is determined as unrealistic.
A.6.5 Planning Horizon
For single cube tasks, the planning horizon is $200$ same to the episode length, and $500$ for others to train the models with enough data.
A.6.6 Value-Learning Policy
@@
Similar to antmaze tasks, we applied value-learning policy (dql) by achieving the subgoals suggested by planners for $10$ steps further planned states.
A.7 Visual Pointmaze
A.7.1 Setup
@@
Because diffusion‐based planners typically operate on state‐action representations, we first learn an image encoder via a Variational Autoencoder (VAE) (vae). The resulting latent states serve as inputs to the diffusion models. Notably, these latent encodings do not include explicit positional information, so the planner must infer underlying geometry and navigate the maze indirectly. We design an inverse dynamics model (a small network of linear layers) to convert consecutive latent observations into actions: $\hat{a}_{t}=\mathrm{InvDyna}\bigl{(}(x_{t-1},\,x_{t}),\,x_{t+1}\bigr{)}$. This replaces the heuristic controller of the original Diffuser, which cannot be directly applied here due to partial observability.
A.7.2 Guidance
@@
To guide the diffusion process, we introduce a position estimator trained on ground‐truth coordinates. Although this estimator provides an approximate distance to the goal, the image encoder is not updated from this learning signal and it does not supply the planner with the full underlying state. Therefore, the image encoder still sees only the pixel‐level data without direct positional labels, thereby preserving partial observability in the core planning procedure. This design choice highlights an open‐ended question of how best to incorporate task‐relevant signals in visual domains. The same guidance function with maze task, Appendix A.5.2 is applied.
A.7.3 Guidance Set
@@
For this task, we applied $[0,0.1,0.5,1,2]$ guidance set for MCTD and MCTD-Replanning.
A.7.4 Dataset Generation and Planning Horizon
@@
We follow the OGBench (4_park2024ogbench) dataset generation script for PointMaze, with trajectory horizons of 1000 for Medium and Large mazes. To enhance training, we use reduced planning horizons (500 for Medium/Large) and apply a sliding-window technique, as in our earlier pointmaze setup.
A.7.5 Variational Autoencoder
@@
We pretrained a VAE (vae) to encode $64\times 64\times 3$ RGB observations into an $8$-dimensional latent space, compressing high-dimensional inputs into a structured representation for downstream planning. It represents the latent variable $z$ as a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$, using the reparameterization trick:
$z=\mu+\sigma\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (5)
We utilized this sampled latent embedding $z$ as the input representation for downstreaming tasks.
A.7.6 Inverse Dynamics Model and Position Estimator
@@
We trained Inverse Dynamics and Position Estimator models using an offline dataset encoded by the pretrained VAE. Both models are implemented as Multi-Layer Perceptrons (MLPs).
@@
Inverse Dynamics. This model predicts the action $\hat{a}_{t}$ required to transition from the current state $x_{t}$ to the next state $x_{t+1}$. To incorporate velocity-related information in a partially observable environment, the model takes both the previous state $x_{t-1}$ and the current state $x_{t}$ as inputs. The action is predicted as follows:
$\hat{a}_{t}=\mathrm{InvDyna}\bigl{(}(x_{t-1},\,x_{t}),\,x_{t+1}\bigr{)}$ (6)
For the initial state, where $x_{t-1}$ is unavailable, we use $x_{t}$ twice as a substitute.
@@
Position Estimator. The Position Estimator predicts the agent’s current position in the latent space encoded by the VAE, providing a weak positional signal to aid planning while preserving partial observability.
A.7.7 Replanning is Necessary to Solve Complex Environments

### v5

@@ A.3 Model Hyperparameters
For reproducibility, we detail the hyperparameters used in our experiments. These settings were selected based on prior work and empirical tuning to ensure stable training and evaluation. Nearly identical hyperparameters were applied consistently across all tasks, except where task-specific configurations were necessary, which are discussed in their respective sections.
@@ A.3.1 Diffuser
Our Diffuser implementation is based on the official repository (https://github.com/jannerm/diffuser). While retaining the core architecture and training procedure, we introduce the following modifications to align with our experimental setup:
  • Guidance Function: Whereas the original Diffuser applied goal-inpainting guidance for point-maze tasks, we employ a distance-based reward function to ensure a fair comparison across all baselines.
  • Replanning Variant: To evaluate the impact of inference-time iterative refinement, we implemented a replanning strategy that periodically re-samples trajectories based on the agent’s updated state.
  • Random Search Variant: To assess the effect of scaling inference-time computation, we implement the “Sample-Score-Rank” variant of Diffuser from (Zhou et al., 2024). Furthermore, to incorporate varied guidance, we assign a guidance scale to each sample, drawn uniformly from the set $\{0.01,0.05,0.1,0.2,0.3\}$.
@@ A.3.2 Diffusion Forcing
Our implementation of Diffusion Forcing is based on the officially released source code (https://github.com/buoyancy99/diffusion-forcing). The following modifications were made:
  • Replanning Frequency: Since Diffusion Forcing natively supports inference-time replanning, we evaluated it under comparable conditions by applying periodic plan refinement, similar to our Diffuser-Replanning variant.
  • Transformer Backbone: To maintain consistency across models, we utilized the Transformer-based backbone provided in the Diffusion Forcing repository.
@@ A.3.3 Monte Carlo Tree Diffusion (MCTD)
Our MCTD model inherits most hyperparameters from Diffusion Forcing. However, we introduce specific modifications and add several new hyperparameters related to the tree search component.
@@ A.3.4 Value-Learning Policy
To address the challenges of complex action spaces in specific tasks, we integrate a value-learning policy adapted from the PlanDQ source code (Chen et al., 2024b) (https://github.com/changchencc/plandq).
@@ A.4 The Evaluation Details
We followed the OGBench (Park et al., 2025) task configurations, 5 tasks per each environment. For each task, we evaluated 10 random seeds per each model, and reported the averaged success rates and their standard deviations.
A.5 Maze Navigation with Point-Mass and Ant Robots
A.6 Modifications from OGBench
Although the original OGBench datasets include random noise in the start and goal positions, we eliminated this noise to isolate model performance from environmental randomness. This modification facilitates clearer comparisons across methods and ensures that performance differences can be attributed to planning capabilities rather than uncontrolled stochasticity. Additionally, we incorporate velocity into the state representation, which the original benchmark does not provide, to apply the heuristic controller for Point-Maze tasks implemented in (Janner et al., 2022). For the Ant-Maze Giant task, we found cases where well-generated plans failed due to inaccurate execution. To prevent this, we increased the episode length from $1000$ to $1500$.
A.7 Guidance Function
To plan long-horizon trajectories while maintaining the effectiveness to reach the goal as quickly as possible, we applied a guidance function to minimize the distance between each state and the goal, represented as $\sum_{i}||x_{i}-g||_{2}$ (Chen et al., 2024a). We note that for a fair comparison, we applied the same guidance style across all baselines. For example, Diffuser originally used goal-inpainting guidance for Point-Maze, but in this evaluation, we used the distance-based guidance instead.
A.8 Guidance Scale
We applied different guidance scales to MCTD for different tasks. For Point-Maze Medium and Large, a scale set of $[0,0.1,0.5,1,2]$ was used. For Point-Maze Giant, we applied $[0.5,1,2,3,4]$. For the Ant-Maze tasks, a set of $[0,1,2,3,4,5]$ was used.
A.9 Planning Horizon
For the Medium and Large mazes, we applied a planning horizon of $500$, and for the Giant maze, a horizon of $1000$. The trajectory horizons in the original datasets are $1000$ for Medium and Large, and $2000$ for Giant. We used a shorter planning horizon than the dataset horizon to generate more data through the sliding window technique.
A.10 Heuristic Controller and Value-Guided Policy
For Point-Maze, we adopted the heuristic controller designed in (Janner et al., 2022). For Ant-Maze, we used a value-guided policy (Wang et al., 2023), as was done in (Chen et al., 2024b). However, unlike (Chen et al., 2024b), we only provide the planned position $10$ steps ahead as the subgoal.
A.11 MCTD Reward Function
We designed a heuristic reward function. One component of this function penalizes large, physically impossible position changes between consecutive states. Another component provides a reward for reaching the goal. Reaching the goal earlier yields a larger reward, calculated as $r=(H-t)/H$, where $H$ is the total horizon length and $t$ is the current timestep.
@@ A.11 MCTD Reward Function
A.11.1 Value-Learning Policy Variance
For the AntMaze tasks, we employed a value-learning policy following (Chen et al., 2024b). We observed instances where plans were generated successfully, but the policy subsequently failed to execute them effectively. To account for this execution variance, we conducted three trials per task and seed combination for all AntMaze experiments, reporting the maximum success rate.
A.11.2 Short-Horizon (Stitch) Results
To examine performance when only short-horizon trajectories are available during training, we evaluated our method on the stitch datasets, with results summarized in Table 12. The trajectory horizon in these datasets is significantly shorter than the number of steps required to reach the goal (e.g., a horizon of 100 steps in the dataset versus over 200 required steps for medium map tasks). Consequently, we applied replanning to Diffuser, Diffusion Forcing, and MCTD. In contrast to the long-horizon dataset experiments, the Diffuser demonstrates superior performance over Diffusion Forcing, suggesting that its implicit stitching capability is more effective. Meanwhile, MCTD outperforms the baselines on the medium-sized map due to its multi-sample exploration. However, its search depth was constrained by the limited length of trajectories in the training data, resulting in suboptimal performance on the large map.
Table 13: Comparison Results with Goal-Inpainting Diffuser. Success rates (%) on pointmaze and antmaze environments with medium, large, and giant mazes for the navigate datasets.
@@ A.11.2 Short-Horizon (Stitch) Results
giant-navigate-v0 $0\pm 0$ $0\pm 0$ $12\pm 10$ $20\pm 0$ $\textbf{94}\pm\textbf{9}$
A.11.3 Comparison with Goal-Inpainting Diffuser
@@
The Diffuser (Janner et al., 2022) typically employs a learned guidance function to bias generated trajectories toward high-return outcomes. For PointMaze tasks, however, an alternative approach—goal-inpainting guidance—was introduced (Janner et al., 2022), which improves single-pass planning by ”imagining” an intermediate trajectory whose final state is the designated goal. In the context of Diffuser-Replanning, we adapt goal-inpainting by providing the experienced trajectory as contextual information rather than as a direct denoising target.
@@
As shown in Table 13, goal-inpainting outperforms the standard Diffuser with heuristic guidance across medium, large, and giant pointmazes. By treating the goal as the terminal state to inpaint, the Diffuser model effectively samples trajectories from both ends, leveraging local convolutional mechanisms to reconcile the start and goal conditions. Although this approach yields marked performance gains, MCTD remains superior in long-horizon scenarios, outperforming both the base and replanning variants of the goal-inpainting Diffuser. The key distinction lies in MCTD’s tree-structured partial denoising and adaptive branching; while goal-inpainting enhances single-pass generation, it lacks the iterative exploration-exploitation loop of an explicit tree search. Moreover, the advantage of two-ended trajectory imagination diminishes as the maze size increases (e.g., giant mazes). Consequently, MCTD uncovers robust solutions even where goal-inpainting shows diminishing returns, underscoring the benefits of tree-based refinement in complex, long-range planning tasks.
A.12 Robot Arm Cube Manipulation
A.12.1 Object-Wise Guidance
@@
Given that the robot arm can only manipulate a single cube at a time, we adopt an object-wise guidance mechanism in MCTD. Rather than applying a single, holistic diffusion schedule to all cubes simultaneously, MCTD treats the selection of which cube to move as a meta-action. This design partially mitigates issues arising from plans that attempt to move multiple cubes concurrently, although multi-object manipulation still requires careful sequencing to avoid suboptimal interleaving.
A.12.2 Guidance Function
@@
The guidance function is the same as that used for the maze tasks (Appendix A.7). For MCTD and MCTD-Replanning, the guidance is applied on an object-specific basis, aiming to reduce the distance between an object’s current state and its corresponding goal state.
A.12.3 Guidance Set
@@
For the cube manipulation tasks, a guidance set of $[1,2,4]$ is applied for each object. For instance, in a task with two cubes, the total guidance set size is $6$, comprising a set of three values for each of the two objects.
A.12.4 Reward Function
To avoid generating unrealistic plans, we augmented the reward function from the maze environment with additional constraints. A plan is deemed unrealistic and assigned a reward of 00 if it violates any of the following rules:
  • No Simultaneous Movements: While minor noise in the plan is permissible, significant concurrent movement of multiple objects is forbidden.
  • Stable Final Placement: A cube’s final position must be stable, either on the floor or on top of another cube, not suspended in mid-air.
  • No Collisions: A cube cannot be placed in a location already occupied by another cube.
  • Clearance for Manipulation: A plan cannot move a cube if another cube is positioned on top of it.
A.12.5 Planning Horizon
The planning horizon was set to $200$ for single-cube tasks, matching the episode length. For tasks involving multiple cubes, the horizon was extended to $500$.
A.12.6 Value-Learning Policy
@@
Similar to the antmaze tasks, we employed a value-learning policy (Wang et al., 2023). This policy is trained to achieve subgoals suggested by the planner, which correspond to planned states $10$ steps into the future.
A.12.7 Handling Redundant Operations in Object-Wise Planning
To accelerate object-wise planning, we integrated common, repetitive operations into the framework. These operations include automatically dropping an object when it reaches its goal and pre-positioning the robot arm before executing the plan for each object. Embedding these routines directly mitigates inconsistencies, such as planning a move beyond the arm’s reach or attempting to grasp an object when the gripper is already occupied.
A.13 Visual Pointmaze
A.13.1 Setup
@@
Since diffusion-based planners typically operate on state-action representations, we first learn an image encoder using a Variational Autoencoder (VAE) (Kingma & Welling, 2014). The resulting latent states serve as the input to our diffusion model. Notably, these latent encodings do not contain explicit positional information; consequently, the planner must infer the underlying geometry to navigate the maze indirectly. We design an inverse dynamics model—a compact Multi-Layer Perceptron (MLP)—to map consecutive latent observations to actions. This model replaces the heuristic controller from the original Diffuser architecture, which is not directly applicable in this partially observable setting.
A.13.2 Guidance
@@
To guide the diffusion process, we introduce a position estimator trained on ground-truth coordinates. Although this estimator provides an approximate distance to the goal, its learning signal is not used to update the image encoder. The estimator does not supply the planner with the full underlying state, thereby ensuring the image encoder processes only pixel-level data. This design preserves partial observability in the core planning procedure and highlights the open question of how best to incorporate task-relevant signals in visual domains. We apply the same guidance function as in the maze task described in Appendix A.7.
A.13.3 Guidance Scales
@@
For this task, we use a set of guidance scales $\lambda\in\{0,0.1,0.5,1,2\}$ for both MCTD and MCTD-Replanning.
A.13.4 Dataset Generation and Planning Horizon
@@
We follow the dataset generation script from OGBench (Park et al., 2025) for the PointMaze environment, setting trajectory horizons to 1000 for the Medium and Large mazes. To generate more training data, we use a reduced planning horizon of 500 for these mazes and employ a sliding-window technique, consistent with our earlier pointmaze experiments.
A.13.5 Variational Autoencoder
@@
We pretrain a VAE (Kingma & Welling, 2014) to encode $64\times 64\times 3$ RGB observations into an $8$-dimensional latent space. This compresses the high-dimensional visual inputs into a structured representation suitable for downstream planning. The VAE models the latent variable $z$ as a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$, using the reparameterization trick:
$z=\mu+\sigma\cdot\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(0,I).$ (5)
We use the sampled latent embedding $z$ as the input representation for all downstream tasks.
A.13.6 Inverse Dynamics and Position Estimator Models
@@
We trained the inverse dynamics and position estimator models on an offline dataset of latent states encoded by the pretrained VAE. Both models are implemented as MLPs.
Inverse Dynamics.
@@
This model, $f_{\text{inv}}$, predicts the action $\hat{a}_{t}$ required to transition from state $z_{t}$ to the subsequent state $z_{t+1}$. To account for velocity in a partially observable context, the model is conditioned on both the previous and current latent states, $z_{t-1}$ and $z_{t}$:
$\hat{a}_{t}=f_{\text{inv}}(z_{t-1},z_{t},z_{t+1}).$ (6)
For the initial timestep ($t=0$), where $z_{t-1}$ is unavailable, we set $z_{-1}=z_{0}$.
Position Estimator.
@@
The position estimator predicts the agent’s $(x,y)$ coordinates from the latent state $z_{t}$. It provides a weak positional signal to aid the planner for guidance while maintaining the overall partial observability of the task.
A.13.7 The Necessity of Replanning in Complex Environments